# Adaptive Rank, Reduced Forgetting: Knowledge Retention in Continual Learning Vision-Language Models with Dynamic Rank-Selective LoRA

### v4

@@ Abstract
We investigate whether the pre-trained knowledge in vision-language models (VLMs), such as CLIP, can be retained—or even enhanced—in continual learning (CL) while incorporating new knowledge from the data stream. Existing CL methods primarily focus on continual downstream adaptation using components isolated from pre-trained model (PTM), increasing inference complexity and limiting improvements to the PTM itself; some also retain knowledge relying on additional reference data, leading to high training costs. To address these limitations, we propose a universal and efficient Continual Learning approach for VLM based on Dynamic Rank-Selective LoRA (CoDyRA), which directly improves the PTMs while preserving the existing knowledge from both pre-training and CL. Through analyses on how LoRA rank and placement impact/regularize the learning and forgetting in CL, we propose CoDyRA that adaptively performs rank-minimized parameter updates in different modules, based on their importance to the current data. This ensures a balance between knowledge acquisition (plasticity) and forgetting mitigation (stability). Our method operates without explicit domain or distribution prediction and does not rely on reference data, enabling seamless task integration while maintaining pre-trained capabilities. Moreover, CoDyRA preserves the original model architecture and deployment pipeline, introducing no additional inference overhead. Extensive experiments demonstrate that our approach enhances representations based on new downstream data while retaining pre-trained knowledge, achieving state-of-the-art results.
@@ 1 Introduction
Figure 1: Illustrations of different approaches for CL in VLMs. (a) Continual full fine-tuning is applied to achieve universal CL updates, often incorporating additional reference data to mitigate the forgetting of pre-trained knowledge (e.g., zero-shot prediction on unseen data, as in [63]). It is computationally expensive and suffers from forgetting. (b) New task-specific components isolated from the PTM are added for CL, e.g., [58, 55]. They require task/domain prediction/gating in inference, increasing inference complexity and restricting the PTM’s improvement potential with restricted usage scenarios. (c) We aim to continually train the model via parameter-and-data-efficient adaptation in a universal manner, without requiring reference data or full fine-tuning as in (a). It enhances the PTM while continually learning from the data stream, surpassing the limitations of (b) on unseen data.
Figure 2: Overview of our proposed method: we propose the dynamic rank-selection LoRA that allows the each pre-trained weight matrices to adaptively insert the necessary ranks for downstream adaptation while retaining pre-trained capabilities. After training on each task, the dynamic ranks of the parameter updates are merged back into the pre-trained weights, without incurring additional inference overhead.
@@
Continual learning (CL) [17, 10, 48, 27, 14] focuses on incrementally learning from new data streams without catastrophic forgetting [37, 34] or the need to retrain with all seen data. With the advances of large-scale pre-trained models (PTMs), recent studies increasingly perform CL with these models [66, 35, 47, 22, 55]. Many PTM-based CL approaches primarily focus on continual adaptation for downstream tasks (e.g., class-incremental learning, CIL [43, 52, 22, 35]), with an emphasis on downstream task performance. While task-specific designs [66, 35, 55] offer promising improvements in adaptation performance, preserving the inherent capabilities of PTMs during continual learning is crucial for maintaining and enhancing generalizable representations. This has garnered growing interest in research areas such as multi-domain continual learning for VLMs [63], or in continual pre-training to enhance PTMs over time [42]. In this paper, we focus on CL of VLMs like CLIP [40] with multi-domain data.
@@
Some approaches, such as ZSCL [63], aim to continually update PTMs by incrementally incorporating new data with full fine-tuning while mitigating forgetting of pre-trained knowledge through memory replay [27, 52, 41] using additional reference data. However, these methods incur high training costs due to the reliance on additional data and extensive retraining (Fig. 1(a)), yet they still struggle with forgetting. To retain PTM capabilities during CL on downstream tasks, some approaches [58, 55] introduce task-specific components that isolated from the PTM during the CL stream, which require task/domain prediction or gating at inference. This increases inference complexity and limits the model’s adaptability to unseen data, restricting its generalization potential (Fig. 1(b)).
@@
In this work, we propose a practical and efficient CL approach for VLMs that continually trains and updates pre-trained models without relying on downstream task-specific components. Our method preserves and even enhances the generalizable representations of PTMs while seamlessly integrating new knowledge from data streams. To achieve efficient and effective learning, we conduct comprehensive analyses of low-rank adaptation (LoRA) [20] for CL in PTMs. Rather than restricting LoRA to fixed locations as in prior works [28, 16], we extend it to all parameter weights for a thorough investigation, aiming to approximate the capabilities of full fine-tuning. We analyze how the rank and placement of LoRA updates impact or regularize the learning dynamics, influencing both downstream adaptation and the retention of existing knowledge (Sec. 3.2).
@@
Our analyses reveal that not all LoRA ranks and placements contribute equally to downstream learning and forgetting. Generally, for arbitrary weights, relatively higher-rank LoRA facilitates learning new tasks but tends to increase forgetting, while lower-rank LoRA mitigates forgetting but limits adaptation. [^1] Similar findings are also observed in the language domain [5]. However, we identify that there usually exists a balance between plasticity (learning new knowledge) and stability (regularizing forgetting) at a moderately small (but not too small) rank, potentially maximizing CL benefits. Moreover, we observe that this balance varies significantly across different parameter placements.
@@
The analysis results motivate us to propose Continual Dynamic Rank-Selective LoRA (CoDyRA), which performs parameter updates via dynamically optimizing for a minimized LoRA rank suitable for the current data. The rank selection for each LoRA is achieved through a Singular Value Decomposition (SVD)-based reformulation [61, 13, 30, 36, 60], with learnable importance weights corresponding to the singular values. With sparsity-promoting regularization, these learnable weights allow the model to dynamically adjust the contribution of each rank during training. This ensures that only the most essential updates (i.e., LoRA with a minimized rank) are integrated into the pre-trained weights, enabling learning from new data while mitigating forgetting. Our contributions are fourfold:
@@
    We investigate a universal and practical CL paradigm for efficiently updating pre-trained VLMs on new data streams while mitigating forgetting, without relying on reference data or additional components.
@@ 2 Related Work
Continual Learning. Continual learning enables models to sequentially learn new tasks while retaining previously acquired knowledge. Experience replay (ER) methods [32, 3, 7, 29, 8, 56, 57] store subsets of past data to refresh the model on prior tasks during new training. Parameter regularization approaches [23, 1, 59, 2, 21] constrain updates to important weights to preserve past knowledge. Meanwhile, dynamic networks [45, 46, 65, 47, 35, 28, 51, 50, 43, 49] adjust their architecture on the fly and balance learning new information while retaining old knowledge.
Multi-Domain Continual Learning of VLMs. Continual learning of VLMs [22, 62] aims to enable VLMs to sequentially learn across diverse domains while retaining their pre-trained generalization capabilities for previously seen tasks. The approach introduced in ZSCL [63] addresses the challenge of preserving zero-shot capabilities while adapting to new tasks, mitigating the risk of catastrophic forgetting. Subsequent works [58, 44, 55] have focused primarily on continual adaptation to downstream tasks while leveraging pre-trained predictions for unseen data. Recently, [55] proposed a more challenging evaluation setting, where the labels from different domains are mixed during testing.
Low-Rank Adaptation. Low-Rank Adaptation (LoRA) [20] has gained popularity as a parameter-efficient method for fine-tuning large pre-trained models. Recent works enhance LoRA by reformulating parameter updates using SVD, focusing on initialization from the SVD of pre-trained weights [60, 36], treating each rank independently [13, 61, 30], or employing a mixture of ranks [53]. These advancements highlight LoRA’s effectiveness in efficiently fine-tuning large models while preserving generalization.
@@ 3.1 Preliminaries
Continual Learning of VLMs. We focus on CL of VLMs with multiple domain data, e.g., multi-domain task-incremental learning (MTIL) [63, 58] and cross-domain task-agnostic incremental learning (X-TAIL) [55]. In the MTIL setting, the model sequentially learns from a set of $T$ tasks. For each task $t$ , the dataset is represented as $\mathcal{D}^{t}=\{({\mathbf{x}}^{t}_{i},y^{t}_{i})\}_{i=1}^{N^{t}}$, where ${\mathbf{x}}^{t}_{i}\in\mathbb{R}^{H\times W\times C}$ is the input image, $y^{t}_{i}\in\mathcal{C}^{t}$ is its corresponding class label, and $N^{t}$ is the number of samples in the task. The category set $\mathcal{C}^{t}=\{y^{t}_{j}\}_{j=1}^{M^{t}}$ represents the class labels for task $t$, consisting of $M^{t}$ classes in total. In the X-TAIL setting, the category set during testing includes classes from both previously seen and unseen domains, denoted as $\mathcal{C}=\mathcal{C}_{\text{seen}}\cup\mathcal{C}_{\text{unseen}}$. Here, $\mathcal{C}_{\text{seen}}=\bigcup_{i=1}^{n}\mathcal{C}_{i}$ includes all classes learned during prior tasks, while $\mathcal{C}_{\text{unseen}}$ represents novel classes that were not encountered during training. This poses a more challenging scenario, as the model must classify test images across a broader range of domains.
@@
Vision-Language Models (VLMs). In vision-language models like CLIP [40], both the image encoder $f_{\theta}$ and text encoder $g_{\psi}$ are implemented as transformers in our case. Given an input image ${\mathbf{x}}\in\mathbb{R}^{H\times W\times C}$, it is segmented into patches and processed by the vision encoder to produce latent visual features ${\mathbf{z}}^{\text{V}}=f_{\theta}({\mathbf{x}})\in\mathbb{R}^{D}$. For text, the class label $y$ is embedded into prompts like “a photo of a [CLS]”, where [CLS] is the class name, resulting in text input ${\mathbf{t}}$. This prompt is encoded by the text encoder, yielding ${\mathbf{z}}^{\text{T}}=g_{\psi}({\mathbf{t}})\in\mathbb{R}^{D}$. During inference, the probability of classifying image ${\mathbf{x}}$ into class $y_{i}\in\{1,\dots,C\}$ is computed using the softmax function as: $p(y_{i}|{\mathbf{x}})=\frac{\exp(\mathtt{sim}({\mathbf{z}}^{\text{V}},{\mathbf{z}}^{\text{T}}{y_{i}})/\tau)}{\sum_{c=1}^{C}\exp(\mathtt{sim}({\mathbf{z}}^{\text{V}},{\mathbf{z}}^{\text{T}}_{y_{c}})/\tau)},$ where $\mathtt{sim}(\cdot)$ is cosine similarity, and $\tau$ is the temperature.
@@ 3.2 Analyses on LoRA Location and Rank
Applying LoRA on all weights. As shown in Fig. 4 and Fig. 4, our analysis of LoRA placement shows that applying LoRA to specific modules alone may limit learning capacity in different ways: (1) The different LoRA placement locations significantly affect downstream performance and the forgetting/retention of pre-trained capabilities; (2) Updating only the vision encoder leads to more forgetting of pre-trained capabilities; (3) Updating only the attention modules helps mitigate forgetting but limits the learning ability on new tasks. These observations align with the distinct roles of different PTM components. Rather than manually selecting LoRA placement, we apply it across all weights and optimize for an adaptive configuration.
@@
Figure 3: Evaluations of downstream task adaptation and zero-shot capability retention after training CLIP with different LoRA insertion locations and ranks. Different shapes represent encoders, colors indicate transformer modules, and sizes denote ranks.
@@ 3.3 Proposed CoDyRA
As shown in Fig. 2, we introduce dynamic rank-selective parameter updates for all pre-trained weight matrices within the Attention and MLP modules of both the vision and text encoders. These updates are adaptively inserted based on the importance of each rank for the current training data. For clarity, during training on task $t$, we denote the pre-trained weights—or those updated from the previous tasks—as $\{\mathbf{W}^{t,m}_{0}\}_{m=1}^{M}$, where $M$ represents the total number of pre-trained weight matrices augmented with LoRA. After training on each task $t$, these dynamic rank updates are merged back into the original pre-trained weights.
@@
To automatically and adaptively decide the rank in the LoRA updates, we introduce a learnable importance weight vector $\mathbf{w}^{t,m}\in\mathbb{R}^{r}$ and incorporate them into the formulation of LoRA as follows:
@@
Table 1: Comparison of different CL methods on X-TAIL for each domain in terms of “Transfer”, “Average”, and “Last” scores (%). The best and the second best results are highlighted in red and blue, respectively. Methods marked with $\dagger$ indicate the use of domain prediction or distribution detection techniques, as in [58, 55].
@@
Continually enhancing the pre-trained model. After training, ranks with zero importance weights are pruned, while non-zero ranks are merged into the pre-trained weights by substituting Eq. (2) into the original LoRA formulation Eq. (1). When a new task is introduced, a fresh set of CoDyRA modules is initialized, and computations involve only these newly initialized CoDyRA modules alongside the continually updated pre-trained model. This approach enables the model to efficiently leverage accumulated knowledge while maintaining a minimal computational footprint, ensuring no additional overhead.
@@
Table 2: Comparison with state-of-the-art methods on 5-shot MTIL setting in terms of “Transfer”, “Average”, and “Last” scores (%). The best and the second best results are highlighted in red and blue, respectively. Methods marked with $\dagger$ indicate the use of domain prediction or distribution detection techniques, as in [58, 55].

### v5

@@ Abstract
Continual learning (CL) aims to accumulate knowledge from sequential data and task streams. Leveraging their strong generalization and flexibility, pre-trained vision-language embedding models such as CLIP (Contrastive Language-Image Pre-training) have been widely adopted and validated in CL. In addition to learning new knowledge, we investigate whether the pre-trained knowledge in CLIP, can be retained or even enhanced, in CL, while incorporating new knowledge from a data stream. Existing CL methods primarily focus on continual downstream adaptation using components isolated from the pre-trained model (PTM), increasing inference complexity and limiting improvements to the PTM itself; some also retain knowledge by relying on additional reference data, resulting in high training costs. To address these limitations, we propose a universal and efficient CL approach for CLIP based on Dynamic Rank-Selective LoRA (CoDyRA), which directly improves the PTMs while preserving the existing knowledge from both pre-training and CL. By analyzing how LoRA rank and placement affect learning and forgetting in CL, we design CoDyRA that adaptively performs rank-minimized parameter updates in different modules, based on their importance to the current data. This ensures a balance between knowledge acquisition (plasticity) and forgetting mitigation (stability). Our method operates without explicit domain or distribution prediction and does not rely on reference data, enabling seamless task integration while maintaining pre-trained capabilities. Moreover, CoDyRA preserves the original model architecture and deployment pipeline, introducing no additional inference overhead. Extensive experiments show that our approach enhances representations for new downstream data while retaining pre-trained knowledge, achieving state-of-the-art results.
@@ 1 Introduction
Continual learning (CL) [18, 9, 49, 29, 13] focuses on incrementally learning from new data streams without catastrophic forgetting [39, 36, 17] or the need to retrain on all previously seen data. With the rise of large-scale pre-trained models (PTMs), recent studies increasingly perform CL in this context [70, 37, 48, 24, 58]. Many PTM-based CL approaches primarily focus on continual adaptation for downstream tasks (e.g., class-incremental learning, CIL [44, 54, 24, 37]), with an emphasis on downstream task performance. While these task-specific designs [70, 37, 58] yield promising results, preserving the inherent capabilities of PTMs during CL is critical to maintaining and improving generalizable representations. Benefiting from the strong generalization and flexibility of jointly learned vision-text embeddings, vision-language models such as CLIP have been widely adopted and studied in CL [24, 66, 50, 69], including continual updating [17] or multi-domain CL [66] of CLIP.
@@
Some approaches, such as ZSCL [66], aim to update PTM, i.e., CLIP, by incrementally incorporating new data through full fine-tuning while mitigating forgetting via memory replay [29, 54, 43] using additional reference data. However, this leads to high training costs and still struggles with forgetting (Fig. 1(a)). Other methods [61, 58] retain PTM capabilities during downstream CL by introducing task-specific components that are isolated from the PTM, requiring task/domain prediction or gating at inference. This increases inference complexity and limits adaptability to unseen data, restricting the model’s generalization potential (Fig. 1(b)).
Figure 1: CL methods for CLIP: (a) Continual full fine-tuning uses reference data to apply universal updates and mitigate forgetting of pre-trained knowledge (e.g., zero-shot performance on unseen data [66]), but is computationally expensive and still susceptible to forgetting; (b) Task-specific modular methods (e.g., [61, 58]) introduce isolated components that require domain prediction or gating at inference, increasing complexity and limiting adaptability and generalization; (c) Our approach enables universal, parameter- and data-efficient adaptation without reference data or task-specific modules. It continually improves the pre-trained model and achieves superior performance on unseen data.
@@
In this work, we propose a practical and efficient CL approach for CLIP that continually trains and updates pre-trained models without relying on downstream task-specific components. Our method preserves and even enhances the generalizable representations of PTMs while seamlessly integrating new knowledge from data streams. To achieve efficient and effective learning, we conduct comprehensive analyses of low-rank adaptation (LoRA) [21] for CL in PTMs. Rather than restricting LoRA to fixed locations as in prior works [30, 16], we extend it to all parameter weights for a thorough investigation, aiming to approximate the capabilities of full fine-tuning. We analyze how the rank and placement of LoRA updates impact or regularize the learning dynamics, influencing both downstream adaptation and the retention of existing knowledge (Sec. 3.2).
@@
Our analyses reveal that not all LoRA ranks and placements contribute equally to downstream learning and forgetting. Generally, for arbitrary weights, relatively higher-rank LoRA facilitates learning new tasks but tends to increase forgetting, while lower-rank LoRA mitigates forgetting but limits adaptation. However, we identify that there usually exists a balance between plasticity (learning new knowledge) and stability (regularizing forgetting) at a moderately small (but not too small) rank, potentially maximizing CL benefits. Moreover, we observe that this balance varies significantly across different parameter placements.
@@
The analysis results motivate us to propose Continual Dynamic Rank-Selective LoRA (CoDyRA), which performs parameter updates by dynamically optimizing for a minimized LoRA rank suitable for the current data. The rank selection for each LoRA is achieved through a Singular Value Decomposition (SVD)-based reformulation [64, 12, 32, 38, 63], with learnable importance weights corresponding to the singular values. With sparsity-promoting regularization, these learnable weights allow the model to dynamically adjust the contribution of each rank during training. This ensures that only the most essential updates (i.e., LoRA with a minimized rank) are integrated into the pre-trained weights, enabling learning from new data while mitigating forgetting. Our contributions are fourfold:
@@
    We investigate a universal and practical CL paradigm for efficiently updating pre-trained CLIP on new data streams while mitigating forgetting, without reference data or additional components.
@@ 2 Related Work
Continual Learning. Continual learning enables models to sequentially learn new tasks while retaining previously acquired knowledge. Experience replay (ER) methods [34, 3, 6, 31, 7, 59, 60] store subsets of past data to refresh the model on prior tasks during new training. Parameter regularization approaches [25, 1, 62, 2, 23] constrain updates to important weights to preserve past knowledge. Meanwhile, dynamic networks [46, 47, 68, 48, 37, 30, 53, 52, 44, 51] adjust their architecture on the fly, balancing the acquisition of new information with the retention of old knowledge.
Continual Learning of CLIP. Continual learning of CLIP [24, 65] aims to enable PTMs to sequentially learn across diverse domains while retaining their pre-trained generalization capabilities for previously seen tasks. [17] continually adapt CLIP to temporal distribution shifts via timestamped data updates. The approach introduced in ZSCL [66] addresses the challenge of preserving zero-shot capabilities while adapting to new tasks, mitigating the risk of catastrophic forgetting. Subsequent works [61, 45, 58] have focused primarily on continual adaptation to downstream tasks while leveraging pre-trained predictions for unseen data. Recently, [58] proposed a more challenging evaluation setting, where the labels from different domains are mixed during testing.
Low-Rank Adaptation. Low-Rank Adaptation (LoRA) [21] has gained popularity as a parameter-efficient method for fine-tuning large pre-trained models. Similar to LoRA, [56] learns a small set of task-specific representation vectors at each layer. Building on LoRA’s framework, recent works enhance LoRA by reformulating parameter updates using SVD, focusing on initialization from the SVD of pre-trained weights [63, 38], treating each rank independently [12, 64, 32], or employing a mixture of ranks [55]. These advancements highlight LoRA’s effectiveness in efficiently fine-tuning large models while preserving generalization.
Figure 2: Overview of CoDyRA: we propose the dynamic rank-selection LoRA that allows the each pre-trained weight matrices to adaptively insert the necessary ranks for downstream adaptation while retaining pre-trained capabilities. After training on each task, the dynamic ranks of the parameter updates are merged back into the pre-trained weights, without incurring additional inference overhead.
@@ 3.1 Preliminaries
Vision-language CLIP model. In CLIP model with joint vision-language embedding [42, 22, 14, 28], both the image encoder $f_{\theta}$ and text encoder $g_{\psi}$ are implemented as transformers in our case. Given an input image ${\mathbf{x}}\in\mathbb{R}^{H\times W\times C}$, it is segmented into patches and processed by the vision encoder to produce latent visual features ${\mathbf{z}}^{\text{V}}=f_{\theta}({\mathbf{x}})\in\mathbb{R}^{D}$. For text, the class label $y$ is embedded into prompts like “a photo of a [CLS]”, where [CLS] is the class name, resulting in text input ${\mathbf{t}}$. This prompt is encoded by the text encoder, yielding ${\mathbf{z}}^{\text{T}}=g_{\psi}({\mathbf{t}})\in\mathbb{R}^{D}$. During inference, the probability of classifying image ${\mathbf{x}}$ into class $y_{i}\in\{1,\dots,C\}$ is computed using the softmax function as: $p(y_{i}|{\mathbf{x}})=\frac{\exp(\mathtt{sim}({\mathbf{z}}^{\text{V}},{\mathbf{z}}^{\text{T}}{y_{i}})/\tau)}{\sum_{c=1}^{C}\exp(\mathtt{sim}({\mathbf{z}}^{\text{V}},{\mathbf{z}}^{\text{T}}_{y_{c}})/\tau)},$ where $\mathtt{sim}(\cdot)$ is cosine similarity, and $\tau$ is the temperature.
CL of CLIP. We focus on CL of CLIP with multiple domain data, e.g., multi-domain task-incremental learning (MTIL) [66, 61] and cross-domain task-agnostic incremental learning (X-TAIL) [58]. In the MTIL setting, the model sequentially learns from a set of $T$ tasks. For each task $t$ , the dataset is represented as $\mathcal{D}^{t}=\{({\mathbf{x}}^{t}_{i},y^{t}_{i})\}_{i=1}^{N^{t}}$, where ${\mathbf{x}}^{t}_{i}\in\mathbb{R}^{H\times W\times C}$ is the input image, $y^{t}_{i}\in\mathcal{C}^{t}$ is its corresponding class label, and $N^{t}$ is the number of samples in the task. The category set $\mathcal{C}^{t}=\{y^{t}_{j}\}_{j=1}^{M^{t}}$ represents the class labels for task $t$, consisting of $M^{t}$ classes in total. In the X-TAIL setting, the category set during testing includes classes from both previously seen and unseen domains, denoted as $\mathcal{C}=\mathcal{C}_{\text{seen}}\cup\mathcal{C}_{\text{unseen}}$. Here, $\mathcal{C}_{\text{seen}}=\bigcup_{i=1}^{n}\mathcal{C}_{i}$ includes all classes learned during prior tasks, while $\mathcal{C}_{\text{unseen}}$ represents novel classes that were not encountered during training. This poses a more challenging scenario, as the model must classify test images across a broader range of domains.
@@ 3.2 Analyses on LoRA Location and Rank
Figure 3: Evaluations of downstream task adaptation and zero-shot capability retention when training CLIP with varying LoRA insertion points and ranks. Shapes indicate encoders, colors denote transformer modules, and sizes reflect rank values.
@@
Applying LoRA on all weights. As shown in Fig. 4 and Fig. 4, applying LoRA to specific modules alone may limit learning capacity in different ways: (1) The different LoRA placement locations significantly affect downstream performance and the forgetting/retention of pre-trained capabilities; (2) Updating only the vision encoder leads to more forgetting of pre-trained capabilities; (3) Updating only the attention modules helps mitigate forgetting but limits the learning ability on new tasks. These observations align with the distinct roles of different PTM components. Rather than manually selecting LoRA placement, we apply it across all weights and optimize for an adaptive configuration.
@@ 3.3 Proposed CoDyRA
We introduce dynamic rank-selective parameter updates for all pre-trained weight matrices within the Attention and MLP modules of both the vision and text encoders, as shown in Fig. 2. These updates are adaptively inserted based on the importance of each rank for the current training data. For clarity, during training on task $t$, we denote the pre-trained weights—or those updated from the previous tasks—as $\{\mathbf{W}^{t,m}_{0}\}_{m=1}^{M}$, where $M$ represents the total number of pre-trained weight matrices augmented with LoRA. After training on each task $t$, these dynamic rank updates are merged back into the original pre-trained weights.
@@
To automatically and adaptively decide the rank in the LoRA updates, we introduce a learnable importance weight vector $\mathbf{w}^{t,m}\in\mathbb{R}^{r}$ and incorporate it into the formulation of LoRA as follows:
@@
Table 1: Comparisons on X-TAIL for each domain in terms of “Transfer”, “Average”, and “Last” scores (%). The best and the second best results are highlighted in red and blue, respectively. Methods marked with $\dagger$ indicate the use of domain prediction or distribution detection [61, 58].
@@
Continually enhancing the pre-trained model. After training, ranks with zero importance weights are pruned, while non-zero ranks are merged into the pre-trained weights by substituting Eq. (2) into the formulation Eq. (1). When a new task is introduced, a fresh set of CoDyRA modules is initialized, and computations involve only these newly initialized CoDyRA modules alongside the continually updated pre-trained model. This approach enables the model to efficiently leverage accumulated knowledge while maintaining a minimal computational footprint, ensuring no additional overhead.
Table 2: Comparisons on 5-shot MTIL setting, evaluated in a similar manner as in Table 1.